An economy with $55000000000 annual GDP grow 10% a year for 10 years continuously ,how much value it is going to have after 10 years of growth?
1
Expert's answer
2015-08-01T00:00:48-0400
We can calculate this using the following formula:
FV = PV * (1+R)^N,
Where
FV – the GDP value in 10 years
PV – GDP present value
R – GDP growth rate
N – number of years
So,
FV = $55000000000 * (1+0.1)^10 = $55000000000 * 2.59 = $142450000000